OpenStudy (anonymous):

A rectangular garden is 5 ft longer than it is wide. Its area is 500 ft2. What are its dimensions?

OpenStudy (anonymous):

w*(w+5)=500 w^2+5w=500

Parth (parthkohli):

Let's put an equation here. First of all, the length is 5 + x, if width is x. So, we know that they get multiplied. \(\Large \color{MidnightBlue}{\Rightarrow x(5 + x) = 500 }\) \(\Large \color{MidnightBlue}{\Rightarrow x^2 + 5x - 500 = 0 }\) Here's when quadratic equations come to play.

OpenStudy (anonymous):

Anywhos, (i'm skipping the working) but the width is 20. So the length is 20+5=25

OpenStudy (anonymous):

Thanks everyone!